# Solving Optimal Experiment Design with Mixed-Integer Convex Methods

### v6

@@ D Computational experiments
In this appendix section, we present additional computational experiments. Table 5 show an extended comparison of the performance between the different solution approaches on the A-Optimal Problem, D-Optimal Problem, A-Fusion Problem and D-Fusion Problem. In Tables 6 and 7, the results of the computations on the GTI-Optimal and GTI-Fusion Problem are shown. The Tables 8, 9 and 10 compare the performance of Boscia using different settings.
@@
$\operatorname{Tr}\left(X^{-1}\right)$ Monotic + Adaptive $\operatorname{Tr}\left(X^{-1}\right)$ Secant $\log\operatorname{Tr}\left(X^{-1}\right)$ Secant
Type Corr. Solved after (s) # inst. % solved Time (s) relative gap # nodes % solved Time (s) relative gap # nodes % solved Time (s) relative gap # nodes

### v7

@@ D Computational experiments
In this appendix section, we present additional computational experiments. Table˜5 show an extended comparison of the performance between the different solution approaches on the A-Optimal Problem, D-Optimal Problem, A-Fusion Problem and D-Fusion Problem. In Tables˜6 and 7, the results of the computations on the GTI-Optimal and GTI-Fusion Problem are shown. The Tables˜8, 9 and 10 compare the performance of Boscia using different settings.
@@
$\operatorname{Tr}\left(X^{-1}\right)$
Monotic + Adaptive
$\operatorname{Tr}\left(X^{-1}\right)$
Secant
$\log\operatorname{Tr}\left(X^{-1}\right)$
Secant
Type
Corr.
Solved
after
inst.
solved
Time (s)
relative
gap
# nodes
solved
Time (s)
relative
gap
# nodes
solved
Time (s)
relative
gap
# nodes